4(s/a -1)(s/b -1)(s/c -1) = r/R , prove the given equation LHS=RHS.

Grade:12th pass

1 Answers

Arun
4 years ago
Dear student
 
LHS can be written as
 
4(s-a) (s – b) ( s – c) / abc = 4 s (s-a) (s – b) ( s – c) / abc * s
 
 = 4 (delta)^2 / abc * s = delta /(s*R)
 
 = r/R